OpenStudy (anonymous):

How to find exact value... of sin (2π)/3

OpenStudy (anonymous):

what do you mean? with what means? calculator, sequence, approximation?

OpenStudy (radar):

\[\sqrt{3}/2\]

OpenStudy (radar):

Refer to the unit circle

OpenStudy (anonymous):

Remember that pi is equal to 180 degrees and 2pi is 360 degrees. So, this one is 360/3 or 120 degrees. This angle is in the 2nd quadrant and the angle of reference (always from the x axis) is 60 degrees. The sin of 60 is the square root of 3 over 2. And, since it's in the 2nd quadrant, it is positive. Remember sin is positive in the 1st and 2nd quadrant and negative in the 3rd and 4th.

OpenStudy (radar):

Good answer Beeb

OpenStudy (anonymous):

stole it though...but your hint made me look it up

OpenStudy (anonymous):

\[what is \sin 2\pi ? \it is 0 do you mean \sin(2\pi/3)? this is the same as \sin \pi/3\] beep and radar gave you good hints

OpenStudy (radar):

Should be helpful to alive